OpenStudy (anonymous):

f(x)=ln3x has a tangent line at point x=1/3 , find point (x;29)

OpenStudy (amistre64):

the instructions are a bit vague

OpenStudy (anonymous):

I don't really know the correct translation . But I'll try to make something else out of it .

OpenStudy (anonymous):

f(x)=ln3x has a tangent line at point x=1/3 , find point on the tangent line , where y=29 .

OpenStudy (anonymous):

I thought about first finding y , which of 0 , then derivative 1/x , then slope , which is 3 and then using formula k=y2-y1/x2-x1 .

OpenStudy (amistre64):

ln(3*1/3) = ln(1) = 0 thats a good thought ... lets see if we can refine it

OpenStudy (amistre64):

ln(3x) derives to: 3/3x = 1/x the slope of the line at x=1/3 is then 3 y = mx -mPx + Py is the equation of a line: siince our m=3; and our (Px,Py) is (1/3, 0) 29 = 3x -3(1/3) + 0 29 = 3x -1 3x = 30, x=10

OpenStudy (amistre64):

the formula you proposed for "k" is just the slope formula; we needed the slope intercept form of a line to describe the tangent line itself

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

youre welcome